train A and B are traveling in the same direction on parallel tracks. Train A IS TRAVELING 60 MILES PER HOUR AND TRAIN B IS TRAVELING AT 70 MILES PER HOUR. TRAIN A PASSES A STATION AT 4:10 A.M. IF TRAIN B PASSES THE SAME STATION AT 4:22 A.M. AT WHAT TIME WILL TRAIN B CATCH UP WITH TRAIN A?
Answer by mananth(16949)

there are two methods of working
..
speed of train A = 60 mph
speed of train B = 70 mph
..
train B passes 12 minutes later than Train A
so Train A has traveled 12/60*60 miles = 12 miles in 12 minutes.
..
Let train B catch up with train A after x miles
time taken = x/70
..
meanwhile train A travels x-12/60
..
(x-12)/60 = x/70
70(x-12)=60x
70x-840=60x
70x-60x=840
10x=840
x= 84 miles
speed of train B = 70 mph
time = 84/70
=1.2 hours OR 72 minutes
Train B will catch up after 72 minutes at 5.34 pm
..
train B covers 10 miles extra per hour.
train b has traveled 12 miles
to catch up it will take 12/10 hours = 1.2 hours = 72 minutes. ( so simple)